A car is initially traveling toward SOUTH at a speed of 11.75 m/s. The brakes are applied and the car comes to a stop over a distance of 38.3 m (since the beginning of the application of the brakes). Take SOUTH as the positive direction.

(a) Find the car's acceleration in m/s2. Pay attention to its sign. Keep 2 decimal places.
Answer #1

using 3rd equation of motion

v^2 = u^2 + 2 a d

0 = 11.75^2 + 2* a * ( 38 3)

a = - 1.8 m/s^2

====

Comment before rate in case any doubt, will reply for sure.. goodluck